# Phonon-mediated spin-polarized superconductivity in altermagnets

### v1

@@ V Superconductivity
where $S_{\mathrm{FS}}$ is the length of the FS, $A_{\mathrm{BZ}}$ is the area of the first Brillouin zone, $N_{\mathrm{samp}}$ is the number of points used to sample the FS, $k_{\|}$ signifies the component of k parallel to the FS, while $k_{\perp}$ is the perpendicular component, and $d\varepsilon/dk^{\prime}_{\perp}$ is the derivative of the electron energy perpendicular to the FS.
@@
From Fig. 3 (a), the interaction has a similar disadvantageous sign as discussed for jellium optical phonons. However, the interaction is now stronger for $\textbf{k}\neq\pm\textbf{k}^{\prime}$ than at $\textbf{k}=\pm\textbf{k}^{\prime}$, such as in the corners of Fig. 3 (a), which enables solutions for the gap. If at some $^{\prime}\neq\textbf{k}$ k we have $\bar{V}_{k_{\|}k^{\prime}_{\|}\sigma}>0$ and $|\bar{V}_{k_{\|}k^{\prime}_{\|}\sigma}|>|\bar{V}_{k_{\|}k_{\|}\sigma}|$, it is possible to have non-trivial solutions of the gap equation if $\Delta_{\textbf{k}\sigma}$ and $\Delta_{\textbf{k}^{\prime}\sigma}$ have opposite signs, effectively canceling the negative sign on the RHS of (9). For that reason, our gap solutions tend to have more sign changes on the FS than the effective electron-electron interaction.
@@
We are interested in the largest eigenvalue of (9), $\lambda_{\mathrm{m}}$, which is related to the critical temperature by [8] $k_{B}T_{c}\approx 1.13\hbar\omega_{P}e^{-1/\lambda_{\mathrm{m}}}$, where $\omega_{P}$ is the largest phonon frequency. We parameterize the FS similarly to Ref. [39], and ensure we use enough points to get convergence of $\lambda_{\mathrm{m}}$ to 3 decimals.
@@
Figure 3: (a) shows $N\bar{V}_{\textbf{k}^{\prime}\textbf{k}\sigma}$ for both spin-up (green and purple) and spin-down (red and blue). The values for k are given in yellow and correspond to the maximum of the gap, while position along the Fermi surface gives $\textbf{k}^{\prime}$. (b) illustrates the normalized superconducting gap in arbitrary units for both spin-up, and spin-down with the same color coding as in (a). In both plots, the chemical potential is set to $\mu=0.1t$.
@@
Figure 4: (a) shows the dimensionless coupling and (b) shows the critical temperature, both as functions of the chemical potential for the two cases of including and excluding NNN hopping.
@@
Fig. 4 (a) shows that when including the NNN hopping, the effective coupling is much stronger for low values of the chemical potential. Including NNN hopping means that more electrons couple to the phonons. This can be understood by studying the interactions in Fig. 3 (a), where for $t_{2}=0$, the interaction is much weaker around the center of the arcs and stronger in the corners. The matrix elements of the BCS equation, (9), are essentially the interaction multiplied by the inverse slope of the band. Comparing the interaction with the inverse slopes in Fig. 2 (b), the inverse slopes are larger at the center of the arcs and weaker towards the edges. However, for $t_{2}\approx 0.21t$, the interaction and the slopes are to a larger degree constant along the arcs, which makes the eigenvalues larger.
@@
The dimensionless coupling, $\lambda_{m}$, decreases as the chemical potential approaches the band crossing at $\varepsilon=J_{\mathrm{sd}}S$. As shown in Fig. 2 (b), the bands become increasingly flat when approaching the crossing, especially along the lines $\bm{\Gamma}\textbf{X}$ and $\bm{\Gamma}\textbf{Y}$ which causes the RHS of (9) to diverge. Combined with the unfavorable sign of the interaction, this makes it increasingly difficult to satisfy the conditions for a gap with a large eigenvalue.
@@
The spin-split FSs rule out even-frequency, spin-singlet, even-momentum pairing for zero-momentum Cooper pairs. The most likely scenario with the type of FS of our model appears to be zero-momentum pairing. This provides the greatest possible phase space for interactions, and unlike other models, there is no overlap of the FSs of spin-up and spin-down [50]. Careful ARPES measurements of the gap amplitude $|\Delta_{\textbf{k}}|$ would yield information on the nodal structure of the gap. The presence of more than two nodes on the FS per spin would be a strong indication that the mechanism described in this letter is at play. Identification of nodes can in principle be used to distinguish zero-momentum spin-triplet pairing from finite momentum pairing spin-singlet, since the latter yields $s$-wave or $d_{x^{2}-y^{2}}$-pairing [51, 50]. As noted in Ref. [50], finite momentum pairing is also predicted to yield a distinct Bogoliubov-de Gennes quasiparticle energy spectrum and a distinct density of states compared to zero-momentum pairing. In our case, with a nodal gap, we predict a V-shaped density of states. We emphasize that these predictions should be directly testable on the compounds $\rm{Rb_{1-\delta}V_{2}Te_{2}O}$ and $\rm{KV_{2}Se_{2}O}$ [47, 46].
@@ Acknowledgements.
Supplemental Material
Phonon-mediated spin-polarized superconductivity in altermagnets
S.VII Introduction
In Section S.VIII, we discuss the symmetries of the Lieb lattice. In Section S.IX, we provide details of how the phonon-spectrum of the Lieb-lattice altermagnet Hamiltonian is computed. In Section S.X, we present details of the derivation of the electron-phonon coupling in this particular system. In Section S.XI, we discuss the details of the effective phonon-mediated electron-electron interaction on the Fermi surface, and how it can promote pairing due to the anisotropies of the effective interaction on the Lieb lattice. Section S.XII gives the values of the fixed parameters we have used throughout. (Parameters that are varied are given in the relevant figure captions).
S.VIII Lattice symmetries
We model the altermagnet as a two-dimensional Lieb lattice as proposed in Ref. [39], a figure of the lattice can be found in Fig. 1 (a) of the main text. The Lieb lattice exhibits the eight symmetries
$\begin{array}[]{cccc}\left(E|0\right),&\left(C_{2z}|0\right),&\left(\sigma_{x}|0\right),&\left(\sigma_{y}|0\right),\\[5.0pt] \left(C_{4z}^{+}|\mathcal{T}\right),&\left(C_{4z}^{-}|\mathcal{T}\right),&\left(\sigma_{xy}|\mathcal{T}\right),&\left(\sigma_{\bar{xy}}|\mathcal{T}\right),\end{array}$ (S1)
where E is the identity operation, $C_{\mathrm{2z}}$ is a twofold rotation, $C_{\mathrm{4z}}^{+}$ and $C_{\mathrm{4z}}^{-}$ are fourfold clockwise and anticlockwise rotations respectively, while $\sigma_{x}$, $\sigma_{y}$, $\sigma_{xy}$ and $\sigma_{\bar{xy}}$ are mirror symmetries about the x-axis, y-axis, the $y=x$ diagonal, and the $y=-x$ diagonal, respectively, and $\mathcal{T}$ is the time-reversal operator, effectively flipping the on-site spins. The notation $\left(A|0\right)$ means that A is a symmetry without performing time-reversal, while $\left(B|\mathcal{T}\right)$ indicates that a combination of B and time-reversal is a symmetry.
S.IX Phonon model
@@
$V(\textbf{r}_{11},\dots,\textbf{r}_{NN_{b}})=V(\textbf{R}_{11},\dots,\textbf{R}_{NN_{b}})+\sum_{i\alpha\mu}\left.\partialderivative{V}{r_{i\alpha}^{\mu}}\right|_{\mathrm{eq}}u^{\mu}_{i\alpha}+\frac{1}{2}\sum_{i\alpha\mu}\sum_{j\beta\nu}\left.\frac{\partial^{2}V}{\partial r_{i\alpha}^{\mu}r_{j\beta}^{\nu}}\right|_{\mathrm{eq}}u^{\mu}_{i\alpha}u^{\nu}_{j\beta},$ (S2)
@@
where $\textbf{r}_{i\alpha}=\textbf{R}_{i\alpha}+\textbf{u}_{i\alpha}$, is the instantaneous position of the lattice atoms, N is the total number of unit cells, $N_{b}$ is the number of atoms in the basis, $\mu,\nu\in\{x,y,z\}$ are spatial directions, and $|_{\mathrm{eq.}}$ indicate to evaluate at equilibrium. The first term is a constant, which we can ignore, and the second term is a force, yielding zero when evaluated in equilibrium, so we are only left with the third term. We define the coefficient of the last term as the force constant matrix (FCM) $\Phi^{\alpha\beta}_{\mu\nu}(\textbf{R}_{i}-\textbf{R}_{j})\equiv\left.\partial^{2}V/\partial r_{i\alpha}^{\mu}\partial r_{j\beta}^{\nu}\right|_{\mathrm{eq.}}$. The FCM is related to the potential of the lattice, hence, it must satisfy the same symmetries as the lattice, namely, (S1), which allows us to find constraints on the matrix elements of the FCM. In addition to these symmetries, the FCM satisfies the conditions [66, 69]
$\sum_{j\beta}$ $\Phi_{\mu\nu}^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)=0,$ (S3)
@@
$\Phi_{\mu\nu}^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)=\sum_{\mu^{\prime}\nu^{\prime}}S^{\bm{\top}}_{\mu\mu^{\prime}}\Phi_{\mu^{\prime}\nu^{\prime}}\left(\mathcal{S}\left(\textbf{R}_{i\alpha}\right)-\mathcal{S}\left(\textbf{R}_{j\beta}\right)\right)S_{\nu^{\prime}\nu},$ (S4)
@@
where $\mathcal{S}$ is a symmetry of the lattice, so $\mathcal{S}\left(\textbf{R}_{i\alpha}\right)$ is the transformed vector of $\textbf{R}_{i\alpha}$ under the symmetry, and $S$ is the matrix representation of this symmetry. We interpret $\Phi^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)$ as a bond between atom $i\alpha$ and atom $j\beta$, and separate the symmetries of these bonds in two sets, $G_{I}$ and $G_{R}$. $G_{I}$ is the set of symmetries that leaves the bond invariant up to a lattice translation, while $G_{R}$ contains the symmetries that reverse the bond by switching the positions of the atoms in the bond, up to a lattice translation. The FCM will then satisfy the conditions
@@
$\Phi^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)=\left\{\begin{array}[]{ll}S^{\bm{\top}}\Phi^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)S&\quad\text{if}\>S\in G_{I},\\ S^{\bm{\top}}\Phi^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)^{\bm{\top}}S&\quad\text{if}\>S\in G_{R}.\par\end{array}\right.$ (S5)
@@
Following Ref. [69], we consider connections up to the third nearest neighbors and explicitly find five distinct bonds, while the rest can be expressed in terms of these by applying (S5).
@@
$\Phi^{11}\left(\left(2a,0\right)\right):\quad$ $G_{I}$ =$=\{E,\sigma_{y}\},\quad$ $G_{R}$ =$=\{C_{2z},\sigma_{x}\}\quad$$\xRightarrow{\quad}$ $\quad\Phi^{11}\left(\left(2a,0\right)\right)=\begin{pmatrix}\eta^{1}_{1}&0\\ 0&\eta^{1}_{2}\end{pmatrix},$
$\Phi^{22}\left(\left(2a,0\right)\right):\quad$ $G_{I}$ =$=\{E,\sigma_{y}\},\quad$ $G_{R}$ =$=\{C_{2z},\sigma_{x}\}\quad$$\xRightarrow{\quad}$ $\quad\Phi^{22}\left(\left(2a,0\right)\right)=\begin{pmatrix}\eta^{2}_{1}&0\\ 0&\eta^{2}_{2}\end{pmatrix},$
$\Phi^{33}\left(\left(2a,0\right)\right):\quad$ $G_{I}$ =$=\{E,\sigma_{y}\},\quad$ $G_{R}$ =$=\{C_{2z},\sigma_{x}\}\quad$$\xRightarrow{\quad}$ $\quad\Phi^{33}\left(\left(2a,0\right)\right)=\begin{pmatrix}\eta^{3}_{1}&0\\ 0&\eta^{3}_{2}\end{pmatrix},$
[…]
$\Phi^{13}\left(\left(0,0\right)\right):\quad$ $G_{I}$ =$=\{E\},\quad$ $G_{R}$ =$=\{\sigma_{xy}\}\quad$$\xRightarrow{\quad}$ $\quad\Phi^{13}\left(\left(0,0\right)\right)=\begin{pmatrix}\rho_{1}&\rho_{2}\\ \rho_{3}&\rho_{1}\end{pmatrix}.$

### v2

@@ V Superconductivity
where $S_{\mathrm{FS}}$ denotes the length of the FS and $A_{\mathrm{BZ}}$ the area of the first Brillouin zone. $N_{\mathrm{samp}}$ is the number of sampling points, we split k into components $(k_{\|},k_{\perp})$ tangent and normal to the FS, with $d\varepsilon/dk^{\prime}_{\perp}$ being the derivative of the electron energy in the normal direction to the FS.
If we momentarily consider a jellium model for phonons [76], the strength of the electron-phonon coupling depends only on the momentum transfer. Still considering spin-polarized superconductivity, the symmetrized coupling would be $\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{\textbf{k}\textbf{k}^{\prime}\sigma}=-|g_{\textbf{k}-\textbf{k}^{\prime}}|^{2}/\hbar\omega_{\textbf{k}-\textbf{k}^{\prime}}+|g_{\textbf{k}+\textbf{k}^{\prime}}|^{2}/\hbar\omega_{\textbf{k}+\textbf{k}^{\prime}}$ on the FS for spin $\sigma$.
Let us first consider a simple model for acoustic phonons, $g_{\textbf{q}}=g\sqrt{|\textbf{q}|}$ and $\omega_{\textbf{q}}=u|\textbf{q}|$. Then, $|g_{\textbf{q}}|^{2}/\hbar\omega_{\textbf{q}}=g^{2}/\hbar u$ is momentum independent, and so $\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{\textbf{k}\textbf{k}^{\prime}\sigma}=0$. There is no phonon-mediated spin-polarized Cooper pairing from acoustic phonons within such a simple continuum model. Next, consider a simple model of Einstein phonons, $g_{\textbf{q}}=g|\textbf{q}|$ and $\omega_{\textbf{q}}=\omega_{E}$. Then, $\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{\textbf{k}\textbf{k}^{\prime}\sigma}=g^{2}(|\textbf{k}+\textbf{k}^{\prime}|^{2}-|\textbf{k}-\textbf{k}^{\prime}|^{2})/\hbar\omega_{E}.$ Imagining an approximately circular FS around the $\bm{\Gamma}$ point, the symmetrized interaction closely resembles $\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{\textbf{k}\textbf{k}^{\prime}\sigma}=V\cos\phi_{\textbf{k}\textbf{k}^{\prime}}$, with $\phi_{\textbf{k}\textbf{k}^{\prime}}$ being the angle between k and $\textbf{k}^{\prime}$, and $V>0$ some positive energy. When solving the eigenvalue problem in (9) for a positive eigenvalue $\lambda$, it is advantageous for the interaction to be attractive when $\textbf{k}^{\prime}$ is close to k, due to the negative sign on the RHS. A coupling of the form $\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{\textbf{k}\textbf{k}^{\prime}\sigma}=V\cos\phi_{\textbf{k}\textbf{k}^{\prime}}$ with $V>0$ prevents non-trivial gap solutions, since the strongest coupling is always at $\textbf{k}=\pm\textbf{k}^{\prime}$, with a disadvantageous sign. This illustrates the key difference between phonon- and magnon-mediated interactions, since for magnons typically $|g_{\textbf{k}-\textbf{k}^{\prime}}|>|g_{\textbf{k}+\textbf{k}^{\prime}}|$ when $\textbf{k}\approx\textbf{k}^{\prime}$ due to the strength of the interaction increasing for long-wavelength magnons [27]. In contrast, for long-wavelength optical phonons, the coupling tends to vanish. Hence, in simple continuum models, optical phonons are also unlikely to yield spin-polarized superconductivity.
The electron-phonon coupling in (6) depends on both the incoming and outgoing momentum. The simple forms of $g_{\textbf{q}}$ and $\omega_{\textbf{q}}$ of the jellium model are replaced by complex phase factors giving a richer momentum structure on a lattice  [^1] The electron transformation coefficients are real and even in momentum and hence only provide quantitative, not qualitative changes of the symmetrized interaction.. Thus, both acoustic and optical phonons contribute to the symmetrized pairing, giving a non-zero symmetrized interaction in Fig. 3 (a).
@@
From Fig. 3 (a), the interaction has a similar disadvantageous sign as discussed for jellium optical phonons. However, the interaction is now stronger for $\textbf{k}\neq\pm\textbf{k}^{\prime}$ than at $\textbf{k}=\pm\textbf{k}^{\prime}$, such as in the corners of Fig. 3 (a), which enables solutions for the gap. If at some $\textbf{k}^{\prime}\neq\textbf{k}$ we have $\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{k_{\|}k^{\prime}_{\|}\sigma}>0$ and $|\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{k_{\|}k^{\prime}_{\|}\sigma}|>|\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{k_{\|}k_{\|}\sigma}|$, it is possible to have non-trivial solutions of the gap equation if $\Delta_{\textbf{k}\sigma}$ and $\Delta_{\textbf{k}^{\prime}\sigma}$ have opposite signs, effectively canceling the negative sign on the RHS of (9). For that reason, our gap solutions tend to have more sign changes on the FS than the effective electron-electron interaction, see Fig. 3.
@@
We are interested in the largest eigenvalue of (9), $\lambda_{\mathrm{m}}$, which is related to the critical temperature by [8]
$k_{B}T_{c}\approx 1.13\hbar\omega_{P}e^{-\textstyle\frac{1}{\lambda_{\mathrm{m}}}},$ (10)
where $\omega_{P}$ is the largest phonon frequency. Due to the exponential dependence of the critical temperature on the dimensionless coupling, the calculated value of the critical temperature will be sensitive to the chosen values of our parameters. Regardless, our aim is not to pinpoint the exact value of the critical temperature, but rather to show, as a proof of concept, the possibility of achieving finite critical temperatures of phonon-mediated spin-polarized superconductivity. We parameterize the FS similarly to Ref. [39], and ensure we use enough sampling points to get convergence of $\lambda_{\mathrm{m}}$ to 3 decimals.
@@
Figure 3: (a) shows $N\mkern 1.0mu\overline{\mkern-1.0muV\mkern-1.0mu}\mkern 1.0mu_{\textbf{k}^{\prime}\textbf{k}\sigma}$ for both spin-up (green and purple) and spin-down (red and blue). The values for k are given in yellow and correspond to the maximum of the gap, while the position along the Fermi surface gives $\textbf{k}^{\prime}$. (b) illustrates the normalized superconducting gap in arbitrary units for both spin-up, and spin-down with the same color coding as in (a). In both plots, the chemical potential is set to $\mu=0.1t$, see Tab. 1 for the other parameters.
@@
Figure 4: (a) shows the dimensionless coupling and (b) shows the critical temperature, both as functions of the chemical potential for the two cases of including and excluding NNN hopping. See Tab. 1 for the parameters used.
@@
Fig. 4 (a) shows that when including the NNN hopping, the dimensionless coupling is much larger for low values of the chemical potential compared to the case $t_{2}=0$. Including NNN hopping means that more electrons couple to the phonons. This can be understood by studying the interactions in Fig. 3 (a), where for $t_{2}=0$, the interaction is much weaker around the center of the arcs and stronger in the corners. The matrix elements of the BCS equation, (9), are essentially the interaction multiplied by the inverse slope of the band. Comparing the interaction with the inverse slopes in Fig. 2 (b), the inverse slopes are larger at the center of the arcs and weaker towards the edges. However, for $t_{2}\approx 0.21t$, the interaction and the slopes are to a larger degree constant along the arcs, which makes the eigenvalues larger.
@@
The dimensionless coupling, $\lambda_{m}$, decreases as the chemical potential approaches the band crossing at $\varepsilon=J_{\mathrm{sd}}S$. As shown in Fig. 2 (b), the bands become increasingly flat when approaching the crossing, especially along the lines $\bm{\Gamma}\textbf{X}$ and $\bm{\Gamma}\textbf{Y}$, which causes the RHS of (9) to diverge. Combined with the unfavorable sign of the interaction, this makes it increasingly difficult to satisfy the conditions for a gap with a large eigenvalue.
@@
Careful ARPES measurements of the gap amplitude $|\Delta_{\textbf{k}}|$ would yield information on the nodal structure of the gap. Combining this with phase-sensitive measurements can corroborate the presence of nodes [17]. Additionally, confirming the presence of nodes using STM based methods has been proposed [84, 85]. If the number of nodes on the FS per spin corresponds to $p$- or $f$-wave gaps, it would be a strong indication that the mechanism described in this paper is at play. Note that nodal gaps are sensitive to impurities, and so an experimental realization likely requires a rather clean system. Identification of nodes can, in principle, be used to distinguish zero-momentum spin-triplet pairing from finite-momentum spin-singlet pairing, since the latter yields $s$-wave or $d$-wave pairing [55, 54]. As noted in Ref. [54], finite momentum pairing is also predicted to yield a distinct Bogoliubov-de Gennes quasiparticle energy spectrum and a distinct density of states compared to zero-momentum pairing. In our case, with a nodal gap, we predict a V-shaped density of states. We emphasize that these predictions should be directly testable on the compounds $\rm{Rb_{1-\delta}V_{2}Te_{2}O}$, $\rm{KV_{2}Se_{2}O}$, and ${\mathrm{La}}_{2}{\mathrm{O}}_{3}{\mathrm{Mn}}_{2}{\mathrm{Se}}_{2}$ [47, 46, 48].
@@ Acknowledgements.
Appendix A Lattice symmetries
We model the altermagnet as a two-dimensional Lieb lattice as proposed in Ref. [39], a figure of the lattice can be found in Fig. 1 (a). The Lieb lattice exhibits the eight symmetries
$\begin{array}[]{cccc}\left(E|0\right),&\left(C_{2z}|0\right),&\left(\sigma_{x}|0\right),&\left(\sigma_{y}|0\right),\\[5.0pt] \left(C_{4z}^{+}|\mathcal{T}\right),&\left(C_{4z}^{-}|\mathcal{T}\right),&\left(\sigma_{xy}|\mathcal{T}\right),&\left(\sigma_{\mkern 1.0mu\overline{\mkern-1.0muxy\mkern-1.0mu}\mkern 1.0mu}|\mathcal{T}\right),\end{array}$ (11)
where E is the identity operation, $C_{\mathrm{2z}}$ is a twofold rotation, $C_{\mathrm{4z}}^{+}$ and $C_{\mathrm{4z}}^{-}$ are fourfold clockwise and anticlockwise rotations respectively, while $\sigma_{x}$, $\sigma_{y}$, $\sigma_{xy}$ and $\sigma_{\mkern 1.0mu\overline{\mkern-1.0muxy\mkern-1.0mu}\mkern 1.0mu}$ are mirror symmetries about the x-axis, y-axis, the $y=x$ diagonal, and the y=xy=-x diagonal, respectively. $\mathcal{T}$ is the time-reversal operator, effectively flipping the on-site spins. The notation $\left(A|0\right)$ means that A is a symmetry without performing time-reversal, while $\left(B|\mathcal{T}\right)$ indicates that a combination of B and time-reversal is a symmetry.
Appendix B Phonon properties from symmetries
@@
$V(\textbf{r}_{11},\dots,\textbf{r}_{NN_{b}})=V(\textbf{R}_{11},\dots,\textbf{R}_{NN_{b}})$ (12)
@@
$+\sum_{i\alpha\mu}\left.\partialderivative{V}{r_{i\alpha}^{\mu}}\right|_{\mathrm{eq}}u^{\mu}_{i\alpha}+\frac{1}{2}\sum_{i\alpha\mu}\sum_{j\beta\nu}\left.\frac{\partial^{2}V}{\partial r_{i\alpha}^{\mu}r_{j\beta}^{\nu}}\right|_{\mathrm{eq}}u^{\mu}_{i\alpha}u^{\nu}_{j\beta},$
@@
where $\textbf{r}_{i\alpha}=\textbf{R}_{i\alpha}+\textbf{u}_{i\alpha}$, is the instantaneous position of the lattice atoms, N and $N_{b}$ denote the number of unit calls and atoms in the basis, respectively. $\mu,\nu\in\{x,y,z\}$ label Cartesian directions, and $|_{\mathrm{eq.}}$ denote evaluation at equilibrium. The first term is a constant, which we can ignore, and the second term is a force, yielding zero when evaluated in equilibrium, so we are only left with the third term. We define the coefficient of the last term as the force constant matrix (FCM) $\Phi^{\alpha\beta}_{\mu\nu}(\textbf{R}_{i}-\textbf{R}_{j})\equiv\left.\partial^{2}V/\partial r_{i\alpha}^{\mu}\partial r_{j\beta}^{\nu}\right|_{\mathrm{eq.}}$. The FCM is related to the potential of the lattice, hence, it must satisfy the same symmetries as the lattice, namely, (11), which allows us to find constraints on the matrix elements of the FCM. In addition to these symmetries, the FCM satisfies the conditions [75, 78]
$\sum_{j\beta}\Phi_{\mu\nu}^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)=0,$ (13)
@@
$\Phi_{\mu\nu}^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)=\sum_{\mu^{\prime}\nu^{\prime}}S^{{\mathrm{T}}}_{\mu\mu^{\prime}}\Phi_{\mu^{\prime}\nu^{\prime}}\left[(\mathcal{S}\left(\textbf{R}_{i\alpha}\right)-\mathcal{S}\left(\textbf{R}_{j\beta}\right)\right]S_{\nu^{\prime}\nu},$ (14)
@@
where $\mathcal{S}$ is a symmetry of the lattice, so $\mathcal{S}\left(\textbf{R}_{i\alpha}\right)$ is the transformed vector of $\textbf{R}_{i\alpha}$ under the symmetry, and $S$ is the matrix representation of this symmetry. We interpret $\Phi^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)$ as a bond between atom $i\alpha$ and atom $j\beta$, and separate the symmetries of these bonds into two sets, $G_{I}$ and $G_{R}$. $G_{I}$ is the set of symmetries that leave the bond invariant up to a lattice translation, while $G_{R}$ contains the symmetries that reverse the bond by switching the positions of the atoms in the bond, up to a lattice translation. The FCM will then satisfy the conditions
@@
$\Phi^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)=\left\{\begin{array}[]{ll}S^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)S&\quad\text{if}\>S\in G_{I},\\ S^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{\alpha\beta}\left(\textbf{R}_{i}-\textbf{R}_{j}\right)^{\raisebox{1.50694pt}{$\scriptstyle T$}}S&\quad\text{if}\>S\in G_{R}.\par\end{array}\right$. (15)
@@
Following Ref. [78], we consider connections up to the third nearest neighbors and explicitly find five distinct bonds, while the rest can be expressed in terms of these by applying (15).
$\Phi^{11}\left(\left(2a,0\right)\right):\quad$ $G_{I}$ $=\{E,\sigma_{y}\},\quad$ $G_{R}$ $=\{C_{2z},\sigma_{x}\}\quad$ $\xRightarrow{\quad}$ $\quad\Phi^{11}\left(\left(2a,0\right)\right)=\begin{pmatrix}\eta^{1}_{1}&0\\ 0&\eta^{1}_{2}\end{pmatrix}$,
$\Phi^{22}\left(\left(2a,0\right)\right):\quad$ $G_{I}$ $=\{E,\sigma_{y}\},\quad$ $G_{R}$ $=\{C_{2z},\sigma_{x}\}\quad$ $\xRightarrow{\quad}$ $\quad\Phi^{22}\left(\left(2a,0\right)\right)=\begin{pmatrix}\eta^{2}_{1}&0\\ 0&\eta^{2}_{2}\end{pmatrix}$,
$\Phi^{33}\left(\left(2a,0\right)\right):\quad$ $G_{I}$ $=\{E,\sigma_{y}\},\quad$ $G_{R}$ $=\{C_{2z},\sigma_{x}\}\quad$ $\xRightarrow{\quad}$ $\quad\Phi^{33}\left(\left(2a,0\right)\right)=\begin{pmatrix}\eta^{3}_{1}&0\\ 0&\eta^{3}_{2}\end{pmatrix}$,
@@
$\Phi^{12}\left(\left(0,0\right)\right):\quad$ $G_{I}$ $=\{E,\sigma_{y}\},\quad$ $G_{R}$ $=\varnothing\quad$ $\xRightarrow{\quad}$ $\quad\Phi^{12}\left(\left(0,0\right)\right)=\begin{pmatrix}\gamma_{1}&0\\ 0&\gamma_{2}\end{pmatrix}$,
$\Phi^{13}\left(\left(0,0\right)\right):\quad$ $G_{I}$ $=\{E\},\quad$ $G_{R}$ $=\{\sigma_{xy}\}\quad$ $\xRightarrow{\quad}$ $\quad\Phi^{13}\left(\left(0,0\right)\right)=\begin{pmatrix}\rho_{1}&\rho_{2}\\ \rho_{3}&\rho_{1}\end{pmatrix}$.